THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 84 — 1 December 2006

http://www.astro.keele.ac.uk/MCnews

Editors: Jacco van Loon and Snežana Stanimirović

Editorial

Dear Colleagues,

It is our pleasure to present the 84^{th} issue of the Magellanic Clouds Newsletter, featuring contributions of a high quality and importance. Several works, mostly based on Spitzer Space Telescope data, relate to the production of interstellar dust; e.g., evidence for the destruction of dust by supernovae (Williams et al.) or the inefficient production by them (Ercolano et al.), and an interstellar dust:gas ratio similar to that in the winds of dusty red giants (Leroy et al.). Intermediate-mass stars are observed to experience prodigious mass loss even in the Magellanic Clouds, as seen in the dust shells of carbon stars (Lagadec et al.) and only slightly higher masses of their remnants compared to the Milky Way (Villaver et al.), whilst lots of iron is seen in the remnants from type Ia supernovae (Borkowski et al.). Also don’t miss the contribution by Reid and Parker, who use a much enlarged sample of planetary nebulae to derive interesting measurements of the kinematic structure of the LMC, as well as some very worthwhile recent conference presentations.

The next issue will be distributed on the 1^{st} of February; the deadline for contributions is the 31^{st} of January.

Looking back on the year 2006, there is no doubt that we have learnt a lot about the Magellanic Clouds, objects and phenomena observed within the Magellanic Clouds, and the physics of stars, gas and galaxies in general. Hopefully the Magellanic Clouds Newsletter has been able to make some small contribution towards synthesizing all of this, and we would like to thank you for your very positive reactions to the resurrection of the newsletter (please keep advertising it amongst your colleagues and students). We wish you a very happy, healthy and Magellanic 2007.

Editorially Yours,

Jacco van Loon and Snežana Stanimirović

Refereed Journal Papers

The Mass Distribution of the Central Stars of Planetary Nebulae in the Large Magellanic Cloud

Eva Villaver^{1}, Letizia Stanghellini^{2} and Richard A. Shaw^{2}

1STScI/ESA, USA 2NOAO, USA

We present the properties of the central stars from a sample of 54 Planetary Nebulae (PNe) observed in the Large Magellanic Cloud (LMC) with the Hubble Space Telescope Imaging Spectrograph (STIS). The Hubble Space Telescope’s spatial resolution allows us to resolve the central star from its nebula (and line-of-sight stars) at the distance of the LMC, eliminating the dependency on photoionization modeling in the determination of the stellar flux. For the PNe in which the central star is detected we obtain the stellar luminosities by directly measuring the stellar fluxes through broad-band imaging and the stellar temperatures through Zanstra analysis. From the position of the central stars in the HR diagram with respect to theoretical evolutionary tracks, we are able to determine reliable core masses for 21 central stars. By including the central star masses determined in this paper to the 16 obtained previously using the same technique (Villaver et al. 2003), we have increased the sample of central star masses in the LMC to 37, for which we find a non-Gaussian mass distribution. The average central star mass for this sample is ⟨m_{CS,LMC}⟩ = 0.65 ± 0.07 M_{⊙}, slightly higher than the one reported in the literature for both white dwarfs and the central stars of PNe in the Galaxy. If significant, this higher average central star mass in the LMC can be understood in terms of a metallicity dependency on mass-loss rates during the Asymptotic Giant Branch, since the LMC has on average half the metallicity compared to the Galaxy. Finally, for the 37 objects analyzed in the LMC, we do not find any significant correlation between the mass of the central star and the morphology of the nebula.

Accepted for publication in Astrophysical Journal Available from astro-ph/0610079

Low-excitation blobs in the Magellanic Clouds

Meynadier, F.^{1,2} and Heydari-Malayeri, M.^{1}

1LERMA – Observatoire de Paris, France 2UPMC Université Paris 6, France

Aims : We study an unknown, or very poorly known, interstellar H ii component in the Magellanic Clouds. This is the first study ever devoted to this class of objects, which we call Low-excitation blobs (LEBs). Methods : We used low-dispersion spectroscopy carried out at ESO to obtain emission line intensities of Hα, Hβ, and [O iii] (4959+5007) for 15 objects in the Large Magellanic Cloud and 14 objects in the Small Magellanic Cloud. Results are displayed in excitation ([O iii]/Hβ ratio) versus Hβ luminosity diagrams. Results : We show the presence of an LEB component in the Magellanic Clouds and study its relationship with the already known class of high-excitation blobs (HEBs). The newly found LEBs are lower excitation counterparts of HEBs and are powered by less massive exciting stars. Further study of LEBs is expected to provide new pieces of information for a better understanding the low mass end of the upper initial mass function in the Magellanic Clouds.

Accepted for publication in A&A Available from astro-ph/0610189

Dust Destruction in Fast Shocks of Core-Collapse Supernova Remnants in the Large Magellanic Cloud

Brian J. Williams^{1}, Kazimierz J. Borkowski^{1}, Stephen P. Reynolds^{1}, William P. Blair^{3}, Parviz Ghavamian^{3}, Sean P. Hendrick^{5}, Knox S. Long^{7}, Sean Points^{6}, John C. Raymond^{2}, Ravi Sankrit^{3,8}, R. Chris Smith^{6} and P. Frank

Winkler^{4}

1North Carolina State University, USA 2Harvard-Smithsonian Center for Astrophysics, USA 3Johns Hopkins University, USA 4Middlebury College, USA 5Millersville University, USA 6NOAO/CTIO, USA 7Space Telescope Science Institute, USA 8UC Berkeley, Space Sciences Laboratory, USA

We report observations with the MIPS instrument aboard the Spitzer Space Telescope (SST) of four supernova remnants (SNRs) believed to be the result of core-collapse SNe: N132D (0525–69.6), N49B (0525–66.0), N23 (0506–68.0), and 0453–68.5. All four of these SNRs were detected in whole at 24 μm and in part at 70 μm. Comparisons with Chandra broadband X-ray images show an association of infrared (IR) emission with the blast wave. We attribute the observed IR emission to dust that has been collisionally heated by electrons and ions in the hot, X-ray emitting plasma, with grain size distributions appropriate for the LMC and the destruction of small grains via sputtering by ions. As with our earlier analysis of Type Ia SNRs, models can reproduce observed 70/24 μm flux ratios only if effects from sputtering are included, destroying small grains. We calculate the mass of dust swept up by the blast wave in these remnants, and we derive a dust-to-gas mass ratio of several times less than the often assumed value of 0.25% for the LMC. We believe that one explanation for this discrepancy could be porous (fluffy) dust grains.

Accepted for publication in Astrophysical Journal Letters Available from astro-ph/0610166

AGB nucleosynthesis in the Large Magellanic Cloud. Detailed abundance analysis of the RV Tauri star MACHO 47.2496.8.

M. Reyniers^{1}, C. Abia^{2}, H. Van Winckel^{1}, T. Lloyd Evans^{3}, L. Decin^{1}, K. Eriksson^{4} and K.R. Pollard^{5}

1Instituut voor Sterrenkunde, Departement Natuurkunde en Sterrenkunde, K.U.Leuven, Celestijnenlaan 200D, 3001 Leuven, Belgium 2Dpto. Física Teórica y del Cosmos, Universidad de Granada, 18 071, Granada, Spain 3SUPA, School of Physics and Astronomy, University of St. Andrews, North Haugh, St. Andrews, Fife KY16 9SS, Scotland, UK 4Department of Astronomy and Space Physics, Uppsala University, Box 515, 75120 Uppsala, Sweden 5Department of Physics and Astronomy, University of Canterbury, Private Bag 4800, Christchurch, New Zealand

Context. Abundance analysis of post-AGB objects as probes of AGB nucleosynthesis. Aims. A detailed photospheric abundance study is performed on the carbon-rich post-AGB candidate MACHO 47.2496.8 in the LMC. Methods. High-resolution, high signal-to-noise ESO VLT-UVES spectra of MACHO 47.2496.8 are analysed by performing detailed spectrum synthesis modelling using state-of-the-art carbon-rich MARCS atmosphere models. Results. The spectrum of MACHO 47.2496.8 is not only dominated by bands of carbon bearing molecules, but also by lines of atomic transitions of s-process elements. The metallicity of [Fe/H] = −1.4 is surprisingly low for a field LMC star. The C/O ratio, however difficult to quantify, is greater than 2, and the s-process enrichment is large: the light s-process elements are enhanced by 1.2 dex compared to iron ([ls/Fe] = +1.2), while for the heavy s-process elements an even stronger enrichment is measured: [hs/Fe] = +2.1. The lead abundance is comparable to the [hs/Fe]. With its low intrinsic metallicity and its luminosity at the low end of the carbon star luminosity function, the star represents likely the final stage of a low initial mass star. Conclusions. The LMC RV Tauri star MACHO 47.2496.8 is highly carbon and s-process enriched, and is most probable a genuine post-C(N-type) AGB star. This is the first detailed abundance analysis of an extragalactic post-AGB star to date.

Accepted for publication in A&A Available from astro-ph/0610240

A New Population of Planetary Nebulae Discovered in the Large Magellanic Cloud (II): Complete PN Catalogue

Warren A. Reid^{1} and Quentin A. Parker^{1,2}

1Macquarie University, Sydney, Australia 2Anglo-Australian Observatory, Australia

This paper presents accurate homogeneous positions, velocities and other pertinent properties for 460 newly discovered and 169 previously known planetary nebulae (PNe) in the central 25 square degrees bar region of the Large Magellanic Cloud (LMC). Candidate emission sources were discovered using a deep, high resolution Hα map of the LMC obtained by median stacking a dozen 2 hour Hα exposures taken with the UK Schmidt Telescope (UKST). Our spectroscopic followup of more than 2,000 compact (i.e. < 20′′) Hα emission candidates uncovered has tripled the number of PNe in this area. All of the 169 previously known PNe within this region have also been independently recovered and included in this paper to create a homogeneous data set. Of the newly discovered PNe, we classify 291 as “true”, 54 as “likely” and 115 as “possible” based on the strength of photometric and spectroscopic evidence. Radial velocities have been measured using both weighted averaging of emission lines and cross-correlation techniques against high quality templates. Based on the median comparison of the two systems, we define a measurement error of ±4 km/s. A new velocity map of the central 25 square degrees of the LMC, based on results from the combined new and previously known PNe, is presented, indicating an averaged heliocentric velocity differential of 65 km/s perpendicular to the line of nodes for the entire PN population across our survey area. Averaged velocities of our PNe and molecular hydrogen (from the literature) across 37′ × 37′ sub areas are compared. The PNe are found to have a higher vertical velocity dispersion than the H i disk to a maximum of 10 times the spread of the H i disk, in keeping with the findings of Meatheringham et al. (1988). In addition, moving out from the main bar, we find that the PNe population follows a plane which is somewhat warped in relation to the H i disk. We estimate the total PN population of the entire LMC system, based on our R-equiv Hα limiting magnitude of ∼ 22, to be 956 ± 141.

Accepted for publication in MNRAS Available from astro-ph/0610152

Dense, Fe-rich Ejecta in Supernova Remnants DEM L238 and DEM L249: A New Class of Type Ia Supernova?

K. J. Borkowski^{1}, S. P. Hendrick^{2} and S. P. Reynolds^{1}

1North Carolina State University, USA 2Millersville University, USA

We present observations of two LMC supernova remnants (SNRs), DEM L238 and DEM L249, with the Chandra and XMM-Newton X-ray satellites. Bright central emission, surrounded by a faint shell, is present in both remnants. The central emission has an entirely thermal spectrum dominated by strong Fe L-shell lines, with the deduced Fe abundance in excess of solar and not consistent with the LMC abundance. This Fe overabundance leads to the conclusion that DEM L238 and DEM L249 are remnants of thermonuclear (Type Ia) explosions. The shell emission originates in gas swept up and heated by the blast wave. A standard Sedov analysis implies about 50 solar masses in both swept-up shells, SNR ages between 10,000 and 15,000 yr, low (< 0.05 cm^{−3}) preshock densities, and subluminous explosions with energies of 3 × 10^{50} ergs. The central Fe-rich supernova ejecta are close to collisional ionization equilibrium. Their presence is unexpected, because standard Type Ia SNR models predict faint ejecta emission with short ionization ages. Both SNRs belong to a previously unrecognized class of Type Ia SNRs characterized by bright interior emission. Denser than expected ejecta and/or a dense circumstellar medium around the progenitors are required to explain the presence of Fe-rich ejecta in these SNRs. Substantial amounts of circumstellar gas are more likely to be present in explosions of more massive Type Ia progenitors. DEM L238, DEM L249, and similar SNRs could be remnants of “prompt” Type Ia explosions with young (∼ 100 Myr old) progenitors.

Accepted for publication in ApJ Available from astro-ph/0608297

The Peculiar Main Sequence of the Double Star Cluster NGC 2011 in

the Large Magellanic Cloud with Hubble Space Telescope WFPC2 Observations

D. A. Gouliermis^{1}, S. Lianou ^{2}, M. Kontizas^{2}, E. Kontizas^{3} and A. Dapergolas^{3}

1Max-Planck-Institut für Astronomie, Heidelberg, Germany 2University of Athens, Dpt of Astrophysics, Astronomy and Mechanics, Athens, Greece 3National Observatory of Athens, Institute of Astronomy, Athens, Greece

We report the serendipitous discovery of a peculiar main sequence in archived Hubble Space Telescope WFPC2 observations of the young star cluster NGC 2011 in the Large Magellanic Cloud. The bright part of this main sequence exhibits a prominent double, fork-like feature, as if it consists of twin main sequences, one of them being redder. The color-magnitude diagram, constructed from the stars found in the only available WFPC2 field of the cluster, is used to distinguish the stars according to their membership to each of these sequences and to study their spatial distribution. We find that there are two well distinguished populations in the sense that the redder main sequence is dominated by stars that belong to the main body of the cluster, while the stars of the bluer main sequence belong to the surrounding region. Providing that NGC 2011 is a verified binary cluster, with the second companion unfortunately not observed, and taking into account the general region where this cluster is located, we discuss the possible scenarios from both star formation, and early dynamical evolution point-of-view that might explain this unique discovery.

Accepted for publication in Astrophysical Journal Letters Available from astro-ph/0610653

On the Progenitor of Supernova 1987A

M. Parthasarathy^{1}, David Branch^{1}, E. Baron^{1} and David J. Jeffery^{1}

1University of Oklahoma, USA

A previously unpublished ultralow-dispersion spectrum of Sanduleak −69 202, the stellar progenitor of SN 1987A, is presented and the uncertain presupernova evolution of Sanduleak −69 202 is discussed.

Accepted for publication in Bull. Astron. Soc. India Available from astro-ph/0611033

A Post-AGB Star in the Small Magellanic Cloud Observed with the Spitzer Infrared Spectrograph

Kathleen E. Kraemer^{1}, G. C. Sloan^{2}, J. Bernard-Salas^{2}, Stephan D. Price^{1}, Michael P. Egan^{1} and P. R. Wood^{3}

1US Air Force Research Laboratory, USA 2Cornell University, USA 3Mount Stromlo Observatory, Australia

We have observed an evolved star with a rare combination of spectral features, MSX SMC 029, in the Small Magellanic Cloud (SMC) using the low-resolution modules of the Infrared Spectrograph on the Spitzer Space Telescope. A cool dust continuum dominates the spectrum of MSX SMC 029. The spectrum also shows both emission from polycyclic aromatic hydrocarbons (PAHs) and absorption at 13.7 μm from C_{2}H_{2}, a juxtaposition seen in only two other sources, AFGL 2688 and IRAS 13416−6243, both post-asymptotic giant branch (AGB) objects. As in these sources, the PAH spectrum has the unusual trait that the peak emission in the 7-9 μm complex lies beyond 8.0 μm. In addition, the 8.6 μm feature has an intensity as strong as the C-C modes which normally peak between 7.7 and 7.9 μm. The relative flux of the feature at 11.3 μm to that at 8 μm suggests that the PAHs in MSX SMC 029 either have a low ionization fraction or are largely unprocessed. The 13-16 μm wavelength region shows strong absorption features similar to those observed in the post-AGB objects AFGL 618 and SMP LMC 11. This broad absorption may arise from the same molecules which have been identified in those sources: C_{2}H_{2}, C_{4}H_{2}, HC_{3}N, and C_{6}H_{6}. The similarities between MSX SMC 029, AFGL 2688, and AFGL 618 lead us to conclude that MSX SMC 029 has evolved off the AGB in only the past few hundred years, making it the third post-AGB object identified in the SMC.

Accepted for publication in Astrophysical Journal Letters

Spitzer spectroscopy of carbon stars in the Small Magellanic Cloud

Eric Lagadec^{1}, Albert A. Zijlstra^{1}, G.C. Sloan^{2}, Mikako Matsuura^{3,1}, Peter Wood^{4}, G.J. Harris^{5}, Jacco Th. van Loon^{6}, J.A.D.L. Blommaert^{7}, S. Hony^{7}, M.A.T. Groenewegen^{7}, M.W. Feast^{8}, P.A. Whitelock^{9,10,11}, J.W. Menzies^{9}

and M.-R. Cioni^{12}

1University of Manchester, UK 2Cornell University, USA 3NAO, Japan 4Australian National University, Australia 5University College London, UK 6Keele University, UK 7K.U. Leuven, Belgium 8CEA Saclay, France 9University of Cape Town, South Africa 10South African Astronomical Observatory, South Africa 11NASSP, University of Cape Town, South Africa 12University of Edinburgh, UK

We present Spitzer Space telescope spectroscopic observations of 14 carbon-rich AGB stars in the Small Magellanic Cloud. SiC dust is seen in most of the carbon-rich stars but it is weak compared to LMC stars. The SiC feature is strong only for stars with significant dust excess, opposite to what is observed for Galactic stars. We argue that in the SMC, SiC forms at lower temperature than graphite dust, whereas the reverse situation occurs in the Galaxy where SiC condenses at higher temperatures and forms first. Dust input into the interstellar medium by AGB stars consists mostly of carbonaceous dust, with little SiC or silicate dust. Only the two coolest stars show a 30-μm band due to MgS dust. We suggest that this is due to the fact that, in the SMC, mass-losing AGB stars generally have low circumstellar (dust) optical depth and therefore effective heating of dust by the central star does not allow temperatures below the 650 K necessary for MgS to exist as a solid. Gas phase C_{2}H_{2} bands are stronger in the SMC than in the LMC or Galaxy. This is attributed to an increasing C/O ratio at low metallicity. We present a colour-colour diagram based on Spitzer IRAC and MIPS colours to discriminate between O- and C-rich stars. We show that AGB stars in the SMC become carbon stars early in the thermal-pulsing AGB evolution, and remain optically visible for ∼ 6 × 10^{5} yr. For the LMC, this lifetime is ∼ 3 × 10^{5} yr. The superwind phase traced with Spitzer lasts for ∼ 10^{4} yr. Spitzer spectra of a K supergiant and a compact H ii region are also given.

Submitted to MNRAS Available from astro-ph/0611071

Dust yields in clumpy SN shells: SN 1987A revisited

Barbara Ercolano^{1}, Mike Barlow^{2} and Ben Sugerman^{3}

1Harvard-Smithsonian CfA, USA 2UCL, UK 3STSci, USA

We present a study of the effects of clumping on the emergent spectral energy distribution (SED) from dusty supernova (SN) shells illuminated by a diffuse radiation source distributed throughout the medium. These models are appropriate for Type II SNe older than a few hundred days, when the energy input is dominated by γ-rays from ^{56}Co decay. The fully 3D radiation transport problem is solved using a Monte Carlo code, mocassin, and we present a set of models aimed at investigating the sensitivity of the SEDs to various clumping parameters. We find that, contrary to the predictions of analytical prescriptions, the combination of an optical and IR observational data set is sufficient to constrain dust masses even in the case where optically thick clumps are present. Using both smoothly varying and clumped grain density distributions, we obtain new estimates for the mass of dust condensed by the Type II SN 1987A by fitting the optical and infrared spectrophotometric data of Wooden et al. (1993) at two epochs (day 615 and day 775). When using amorphous carbon grains, our best fits to the observational data imply that about 2.0·10^{−4} M_{⊙} of dust had condensed in the envelope of SN 1987A by day 615 and between 2.0·10^{−4} and 4.2·10^{−3} M_{⊙} by day 775. We find that the absence of a silicate emission or absorption feature in the observed mid-IR spectra implies that no more than 15% of the dust formed around SN 1987A can have been in the form of silicate particles. Our models require larger dust masses for the case of graphite grains, namely between 4.2·10^{−4} and 6.6·10^{−4} M_{⊙} at day 615 and between 4.5·10^{−4} and 7.5·10^{−4} M_{⊙} at day 775. From our numerical models we derive dust masses for SN 1987A that are comparable to previous analytic clumped graphite grain mass estimates, and at least two orders of magnitude below the 0.1–0.3 M_{⊙} that have been predicted to condense as dust grains in primordial core collapse supernova ejecta. This low condensation efficiency for SN 1987A is in contrast to the case of SN 2003gd, for which a dust condensation efficiency as large as 0.12 has recently been estimated.

Accepted for publication in MNRAS Available from astro-ph/0611719

The Spitzer Survey of the Small Magellanic Cloud: FIR Emission and Cold Gas in the SMC

Adam Leroy^{1,2}, Alberto Bolatto^{2}, Snežana Stanimirović^{3,2}, Norikazu Mizuno^{4}, Frank Israel^{5} and Caroline Bot^{6}

1MPIA, Germany 2UC Berkeley, USA 3University of Wisconsin, USA 4Nagoya University, Japan 5Sterrewacht Leiden, The Netherlands 6Spitzer Science Center, USA

We present new far infrared maps of the Small Magellanic Cloud (SMC) at 24, 70, and 160 μm obtained as part of the Spitzer Survey of the Small Magellanic Cloud (S^{3}MC, Bolatto et al. 2006). These maps cover most of the active star formation in the SMC Bar and the more quiescent Wing. We combine our maps with literature data to derive the dust surface density across the SMC. We find a total dust mass of M_{dust} = 3 × 10^{5} M_{⊙}, implying a dust-to-hydrogen ratio over the region studied of log_{10} D/H = −2.86, or 1-to-700, which includes H_{2}. Assuming the dust to trace the total gas column, we derive H_{2} surface densities across the SMC. We find a total H_{2} mass M_{H2} = 3.2 × 10^{7} M_{⊙} in a distribution similar to that of the CO, but more extended. We compare profiles of CO and H_{2} around six molecular peaks and find that on average H_{2} is more extended than CO by a factor of ∼ 1.3. The implied CO-to-H_{2} conversion factor over the whole SMC is X_{CO} = 13 ± 1 × 10^{21} cm^{−2} (K km s^{−1})^{−1}. Over the volume occupied by CO we find a lower conversion factor, X_{CO} = 6 ± 1 × 10^{21} cm^{−2} (K km s^{−1})^{−1}, which is still a few times larger than that found using virial mass methods. The molecular peaks have H_{2} surface densities Σ_{H2} ≈ 180 ± 30 M_{⊙} pc^{−2}, similar to those in Milky Way GMCs, and correspondingly low extinctions, A_{V} ∼ 1 – 2 mag. To reconcile these measurements with predictions by the theory of photoionization-regulated star formation, which requires A_{V} ∼ 6, the GMCs must be ∼ 3 times smaller than our 46 pc resolution element. We find that for a given hydrostatic gas pressure, the SMC has a 2 – 3 times lower ratio of molecular to atomic gas than spiral galaxies. Combined with the lower mean densities in the SMC this may explain why this galaxy has only 10% of its gas in the molecular phase.

Accepted for publication in The Astrophysical Journal Available from astro-ph/0611687 and from http://celestial.berkeley.edu/spitzer/publications/

Physical parameters of 15 intermediate-age LMC clusters from modelling of HST colour-magnitude diagrams

Leandro Kerber^{1}, Basílio Santiago^{2} and Enzo Brocato^{3}

1IAG/USP, Sao Paulo, Brazil 2IF/UFRGS, Porto Alegre, Brazil 3INAF - Osservatorio Astronomico di Collurania, Teramo, Italy

We analyzed HST/WFPC2 colour-magnitude diagrams (CMDs) of 15 populous Large Magellanic Cloud (LMC) stellar clusters with ages between ∼0.3 Gyr and ∼3 Gyr. These (V, B − V) CMDs are photometrically homogeneous and typically reach V ∼ 22. Accurate and self-consistent physical parameters (age, metallicity, distance modulus and reddening) were extracted for each cluster by comparing the observed CMDs with synthetic ones. These determinations involved simultaneous statistical comparisons of the main-sequence fiducial line and the red clump position, offering objective and robust criteria to determine the best models. The models explored a regular grid in the parameter space covered by previous results found in the literature. Control experiments were used to test our approach and to quantify formal uncertainties. In general, the best models show a satisfactory fit to the data, constraining well the physical parameters of each cluster. The age-metallicity relation derived by us presents a lower spread than similar results found in the literature for the same clusters. Our results are in accordance with the published ages for the oldest clusters, but reveal a possible underestimation of ages by previous authors for the youngest clusters. Our metallicity results in general agree with the ones based on spectroscopy of giant stars and with recent works involving CMD analyses. The derived distance moduli implied by the most reliable solutions, correlate with the reddening values, as expected from the non-negligible three-dimensional distribution of the clusters within the LMC. The inferred spatial distribution for these clusters is roughly aligned with the LMC disk, being also more scattered than recent numerical predictions, indicating that they were not formed in the LMC disk. The set of ages and metallicities homogeneously derived here can be used to calibrate integrated light studies applied to distant galaxies.

Accepted for publication in Astronomy & Astrophysics Available from astro-ph/0610722

Conference Papers

Star formation history across the Magellanic Clouds and other Local Group galaxies

M.-R.L. Cioni^{1}

1SUPA, School of Physics, University of Edinburgh, Institute for Astronomy, Blackford Hill, Edinburgh EH9 3HJ, United Kingdom

The Ks magnitude distribution of asymptotic giant branch stars provides important constraints on the spatial variation of metallicity and mean-age of the stellar population of galaxies. Here, I present results of the investigation of these parameters across the Magellanic Clouds, M33, NGC 6822 and SagDIG and I discuss the tremendous improvement that new wide-field near-infrared instruments (i.e. VISTA) will provide on our understanding of the global instrinsic as well as dynamical history of these systems.

Oral contribution, published in ”Why galaxies care about AGB stars” (workshop, Vienna, August 7-11, 2006 Available from astro-ph/0611389

Analysis of the eclipsing binaries in the LMC discovered by OGLE: Period distribution and frequency of the short-period binaries

Tsevi Mazeh^{1}, Omer Tamuz^{1} and Pierre North^{2}

1School of Physics and Astronomy, Raymond and Beverly Sackler Faculty of Exact Sciences, Israel 2Laboratoire d’Astrophysique, Ecole Polytechnique Fédérale de Lausanne, Switzerland

We review the results of our analysis of the OGLE LMC eclipsing binaries (Mazeh, Tamuz & North 2006), using EBAS – Eclipsing Binary Automated Solver, an automated algorithm to fit lightcurves of eclipsing binaries (Tamuz, Mazeh & North 2006). After being corrected for observational selection effects, the set of detected eclipsing binaries yielded the period distribution and the frequency of all LMC short-period binaries, and not just the eclipsing systems. Somewhat surprisingly, the period distribution is consistent with a flat distribution in log P between 2 and 10 days. The total number of binaries with periods shorter than 10 days in the LMC was estimated to be about 5000. This figure led us to suggest that (0.7 ± 0.4)% of the main-sequence A- and B-type stars are found in binaries with periods shorter than 10 days. This frequency is substantially smaller than the fraction of binaries found by small Galactic radial-velocity surveys of B stars.

Oral contribution, published in IAU Symp. 240 Available from astro-ph/0611482

Shells in the Magellanic System

Snežana Stanimirović^{1}

1Department of Astronomy, University of Wisconsin, 475 North Charter Street, Madison, WI 53706, USA

The Magellanic System harbors > 800 expanding shells of neutral hydrogen, providing a unique opportunity for statistical investigations. Most of these shells are surprisingly young, 2–10 Myr old, and correlate poorly with young stellar populations. I summarize what we have learned about shell properties and particularly focus on the puzzling correlation between the shell radius and expansion velocity. In the framework of the standard, adiabatic model for shell evolution this tight correlation suggests a coherent burst of star formation across the whole Magellanic System. However, more than one mechanism for shell formation may be taking place.

Oral contribution, published in IAU Symposium 237, ”Triggered Star Formation in a Turbulent ISM” Available from astro-ph/0610387

Reconstructing the Initial Relaxation Time of Young Star Clusters in the Large Magellanic Cloud

Simon Portegies Zwart^{1} and Hui-Chen Chen^{1,2}

1University of Amsterdam, the Netherlands 2NC University, Taiwan

We reconstruct the initial two-body relaxation time at the half mass radius for a sample of young < ∼300 Myr star clusters in the large Magellanic cloud. We achieve this by simulating star clusters with 12288 to 131072 stars using direct N-body integration. The equations of motion of all stars are calculated with high precision direct N-body simulations which include the effects of the evolution of single stars and binaries. We find that the initial relaxation times of the sample of observed clusters in the large Magellanic cloud ranges from about 200 Myr to about 2 Gyr. The reconstructed initial half-mass relaxation times for these clusters has a much narrower distribution than the currently observed distribution, which ranges over more than two orders of magnitude.

Oral contribution, published in “Mass loss from stars and the evolution of stellar clusters”. Proc. of a workshop held in honour of H.J.G.L.M. Lamers, Lunteren, The Netherlands. Eds. A. de Koter, L. Smith and R. Waters (San Francisco: ASP) Available from astro-ph/0610659